OpenStudy (anonymous):

√20x^13y^5 ------------ 5xy^7

OpenStudy (anonymous):

(he √ extends over the y^5, too

OpenStudy (anonymous):

@twicker ?

OpenStudy (anonymous):

√(20/5) = 2 √x^(13-1) = √x^12= √x^4.√x^3 = (x^2)(√x^3) √y^(5--7) = √y^12= √y^4.√y^3 = (y^2)(√y^3) = 2(xy)^2.√((xy)^3))

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf \sqrt{ \frac{ 20x^{13}y^5 }{ 5xy^7 } } }\) try to cancel some exponents first.

OpenStudy (anonymous):

so then it becomes √20x^12/√5y^2?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

i dont get it @twicker

OpenStudy (anonymous):

First step is to split the equation. to values that are possible to find the squareroot easily So, the first part if have taken 20/5 = 4. Then √4 is 2. Set that aside and go to the next part which is the x √x^(13-1) = √x^12= √x^4.√x^3 = (x^2)(√x^3) The last part √y^(5--7) = √y^12= √y^4.√y^3 = (y^2)(√y^3) Combine all now: 2.(x^2)(√x^3).(y^2)(√y^3) =2(xy)^2.√((xy)^3)) read: 2 multiplies by the square of xy multiplied by the squareroot of the cube of xy

OpenStudy (anonymous):

can you draw it out?

OpenStudy (anonymous):

Is this the answer correct because if it is not, there's no point repeating the wrong answer. But if it right, I'll try drawing it out

OpenStudy (anonymous):

i just cant picture the 2(xy)^2.√((xy))^3

OpenStudy (anonymous):

@twicker

OpenStudy (anonymous):

a^x . b^x = (ab)^x eg: 2^2 * 3^2 = 4*9= 36= (2*3)^2= 36